Average Rate of Return

Determine the average rate of return for a project that is estimated to yield total income of $571,500 over six years, has a cost of $685,800, and has a $76,200 residual value. Round to the nearest whole number.
%
Answer #1

Total income of six years = $571,500

Average annual income = Total income of six years/6

= 571,500/6

= $95,250

Average investment = ( Cost of project + Residual value)/2

= (685,800+76,200)/2

= $381,000

Average rate of return = Average annual income / Average investment

= 95,250/381,000

= 25%

Kindly comment if you need further assistance.

Thanks‼!